# Error thresholds of toric codes with transversal logical gates

### v2

@@ C.1 Fold-transversal Hadamard gate
On a 2D torus with equal distance in both directions, the toric code has a transversal Hadamard (tHH) gate [53]. Physically, it can be implemented by first applying $H$ gate to every physical qubit, and then swapping the qubits in a reflection-symmetric way with respect to a diagonal line, see Fig. 12. For toric code on a torus with equal periodicity in both directions, the tHH gate simultaneously performs $H$ gates on the two encoded logical qubits.
@@
Based on the stat-mech model, we now discuss the impact of the tHH gate on the error threshold. For simplicity, we consider a Pauli error channel where $X$ and $Z$ errors occur independently with probabilities $P_{X}$ and $P_{Z}$, in which case $(p_{x},p_{y},p_{z})=((1-P_{Z})P_{X},P_{X}P_{Z},(1-P_{X})P_{Z})$. Therefore, $s^{x}$ and $s^{z}$ are independent variables and we have $J_{y}=0$ in the stat-mech model. Hence, the two flavors of Ising spins, $\{\sigma^{x}\}$ and $\{\sigma^{z}\}$, completely decouple except for the terms involving $s^{x,z}(T+\frac{1}{2})$ due to the tHH gates. We can further exchange all the spins $\sigma^{x}_{l}(t)\leftrightarrow\sigma^{z}_{\bar{l}}(t)$ for all $t\leq T$, after which the two spin flavors are completely decoupled. Since the disorder distributions of $s^{x,z}_{l}$ and $r^{x,z}_{p}$ are translation invariant, we can relabel their spatial locations $l$ and $p$ as $\bar{l}$ and $\bar{p}$ as well. Therefore, the only difference here, compared to the original stat-mech model in Eq. (C), is that the coupling constants $J_{x}$ and $J_{z}$, as well as $K_{x}$ and $K_{z}$, are exchanged for all the terms that involve the terms that are located within $t\leq T$. Consequently, both flavors of spins will locally confine once the error rates exceed the error threshold of quantum memory.
@@ C.2 Fold-transversal $S$ gate
Additionally, the toric code has a transversal $S$ (tSS) gate [53]. Physically, it can be implemented by first applying $S$ or $S^{\dagger}$ gates to every physical qubit on a diagonal, while also applying controlled-$Z$ (CZCZ) gates between pairs of qubit that are reflection-symmetric with respect to the diagonal line, see Fig. 12. For toric code on a torus with equal periodicity in both directions, the tSS gate simultaneously performs $S$ gates on the two encoded logical qubits.
@@
We now discuss the impact of the tSS gate on the error threshold using the stat-mech model. Again, we assume the independent error model where $X$ and $Z$ errors occur independently, so that $J_{z}=0$ in the Hamiltonian. Therefore, the two flavors of Ising spins only couple via the terms that involve $s^{z}_{l}(T+\frac{1}{2})$. Since the disorder distributions are translation invariant, we can mirror-reflect the spatial location of all the $\sigma^{x}$ spins, so that the coupling are geometrically local. We recognize that, the stat-mech model in this case is essentially the same as Eq. (IV) for two toric code blocks undergoing a tCNOT gate, where the two spins for the control and target code blocks become $\{\sigma^{x}\}$ and $\{\sigma^{z}\}$ spins for $X$ and $Z$ errors, and the bit-flip error rates of the two code blocks become the bit-flip and phase-flip error rate of the single code block. Therefore, for the tSS gate, the error threshold of logical $Z$ error will be locally lowered near $t=T$, similar to the target block of the tCNOT gate. In the stat-mech model, this effect can be effectively captured by the increase in local temperature of the $\sigma^{z}$ spins, as sketched in Fig. 10.
@@ Appendix D Stat-mech mapping for generic CSS codes under transversal Clifford gates
In this section, we generalize the stat-mech mapping approach to generic Calderbank-Steane-Shor (CSS) codes that supports global transversal Clifford gates.
@@ D.1 Transversal logical circuit in the binary matrix formalism
The error-detections mechanism of a CSS code under repeated noisy syndrome extractions can also be understood using an extend version the binary matrix formalism. Here, Pauli and syndrome errors over spacetime can be parameterized by a single binary vector $\mathfrak{e}$ over time [^4] In this section, the Fraktur alphabet always denotes the binary vector representation of a spacetime error, while the vector notation is used for binary vectors of length $n$, i.e. the physical errors.:
@@
Here $r_{x/z}(t)\in\{0,1\}^{n_{x/z}}$ represents the absence or existence of the syndrome extraction error at each $X/Z$ check at time $t$. The vertical lines in $\mathfrak{H}$ separate the columns that corresponds to the physical Pauli error and syndrome extraction errors, and the convention will be used in subsequent discussions. Hence, the spacetime detectors formed by repeated syndrome measurements, whose parities are represented by a binary vector $\mathfrak{d}$, can be written in the following matrix form:
@@
where $\mathfrak{d}$ collects the parity of the spacetime detectors, $\mathbbm{1}_{n_{x/z}}$ is the identity matrix of dimension $n_{x/z}$, and $d^{x/z}(t+\frac{1}{2})\in\{0,1\}^{n_{x/z}}$ represents the parity of the $X/Z$ detectors. We will refer to $\mathfrak{H}$ as the spacetime parity check matrix.
@@
$U\Omega U^{T}=\Omega,\ \Omega\equiv\left(\begin{array}[]{cc}\mathbbm{1}_{n}&0\\ 0&\mathbbm{1}_{n}\end{array}\right).$ (53)
@@
where $V_{U}\in SL(n_{x}+n_{z},\mathbb{F}_{2})$ represents the change each individual stabilizer under the transversal gate. In the detector error model, $V$ gives rise to the form of the detectors across the transversal gate. Consider a transversal gate $U$ applied between $t=T$ and $t=T+1$ rounds of syndrome extraction. Assuming the physical Pauli errors do not persist across the transversal gate $U$, the spacetime parity check matrix can be written in the following block form:
@@
$\mathfrak{v}_{U}\mathfrak{H}|_{\bar{U}(T+\frac{1}{2})}=\mathfrak{H}\mathfrak{W}|_{\bar{U}(T+\frac{1}{2})},\ \mathfrak{v}_{U}=\text{diag}\left(\cdots,\mathbbm{1}_{n_{x}+n_{z}},V_{U},V_{U},\cdots\right),\ \mathfrak{W}|_{\bar{U}(T+\frac{1}{2})}=\text{diag}\left(\cdots,{\color[rgb]{0,0,1}\mathbbm{1}_{n_{x}+n_{z}}},U,{\color[rgb]{0,0,1}V_{U}},U,{\color[rgb]{0,0,1}V_{U}},\cdots\right).$ (56)
Here $\mathfrak{v}_{U}$ is a block-diagonal matrix with $V_{U}$ starting from the row block that represents the detectors between $t=T$ and $t=T+1$ (i.e. the row where $V_{U}^{-1}$ is at in Eq. (55)), and $\mathbbm{1}_{n_{x}+n_{z}}$ before that. $\mathfrak{W}|_{\bar{U}(T+\frac{1}{2})}$ is a block-diagonal matrix with alternating $U$ and $V_{U}$ blocks starting from the columns that represent the physical errors at $t=T+\frac{1}{2}$ (i.e. the column on the right of $V_{U}^{-1}$ in Eq. (55)).
@@
$\mathfrak{W}|_{\bar{U}(T+\frac{1}{2})\bar{U}^{\prime}(T^{\prime}+\frac{1}{2})}$ $=\text{diag}\left(\cdots,{\color[rgb]{0,0,1}\mathbbm{1}_{n_{x}+n_{z}}},U,{\color[rgb]{0,0,1}V_{U}},\cdots,UU^{\prime},{\color[rgb]{0,0,1}V_{U}V_{U^{\prime}}},\cdots\right)$
$=\text{diag}\left(\cdots,{\color[rgb]{0,0,1}\mathbbm{1}_{n_{x}+n_{z}}},U,{\color[rgb]{0,0,1}V_{U}},\cdots\right)\text{diag}\left(\cdots,{\color[rgb]{0,0,1}\mathbbm{1}_{n_{x}+n_{z}}},U^{\prime},{\color[rgb]{0,0,1}V_{U^{\prime}}},\cdots\right)=\mathfrak{W}|_{\bar{U}(T+\frac{1}{2})}\mathfrak{W}|_{\bar{U}^{\prime}(T^{\prime}+\frac{1}{2})}.$ (57)
@@ D.2 Stat-mech mapping via the binary matrix formalism
Using the binary matrix formalism, we can now discuss the stat-mech mapping with transversal Clifford gate in the most general setting. Starting from the case of quantum memory. An error model in this case can be specified via the probability distribution
@@
Two spacetime errors $\mathfrak{e}$ and $\mathfrak{e}^{\prime}=\mathfrak{e}+\delta\mathfrak{e}$ trigger the same set of detectors if they differ by a cycle $\delta\mathfrak{e}\in\ker(\mathfrak{H})$. The trivial cycles which do not lead to logical errors are generated by the following four sets of bitstrings
@@
where $l_{z,x}$ are binary vectors that represents logical operators of the original CSS code [^5] It does not matter which time the logical error $l_{z,x}$ is located in $\mathfrak{l}_{x,z}$. In fact, using linear combinations of $\mathfrak{r}^{x,z}$, we can always shift it in time.. In this way, every cycle $\delta\mathfrak{e}$ can be parameterized in the following way:
@@
Putting everything together and using the definition in Eq. (II.1), the logical error rate of the decoder is
@@
The quenched disorder, $\mathfrak{e}$, follows the distribution $\mathrm{prob}(\mathfrak{e})$. In the case of structured CSS codes (such as topological codes) and distribution of errors $\mathrm{prob}(\mathfrak{e})$, the partition function can be brought to the Boltzmann form of a known Hamiltonian.
Now consider the case with transversal logical gate. From Eq. (56), we see that every trivial cycle $\mathfrak{c}^{\prime}$ of the new spacetime parity check matrix $\mathfrak{H}^{\prime}$ with transversal gates is related to the kernel of $\mathfrak{H}$ via $\mathfrak{c}^{\prime}=\mathfrak{W}^{-1}\mathfrak{c}$, $\mathfrak{c}\in\ker(\mathfrak{H})$, where $\mathfrak{W}$ can be computed using Eq. (D.1). Therefore, the new partition function reads
@@
Nevertheless, the two partition functions only differ locally around $t=T$. In fact, away from the transversal gate (i.e. when $t\geq T+1$), it is easy to verify that the new cycles $\mathfrak{W}^{-1}\mathfrak{r}^{x,z}_{i}(t\geq T)$ are linear combinations of the old ones:

### v3

@@ C.1 Fold-transversal Hadamard gate
On a 2D torus with equal distance in both directions, the toric code has a transversal Hadamard (tHH) gate [53]. Physically, it can be implemented by first applying the $H$ gate to every physical qubit, and then swapping the qubits in a reflection-symmetric way with respect to a diagonal line, see Fig. 12. For toric code on a torus with equal periodicity in both directions, the tHH gate simultaneously performs $H$ gates on the two encoded logical qubits.
@@
Based on the stat-mech model, we now discuss the impact of the tHH gate on the error threshold. For simplicity, we consider a Pauli error channel where $X$ and $Z$ errors occur independently with probabilities $P_{X}$ and $P_{Z}$, in which case $(p_{x},p_{y},p_{z})=((1-P_{Z})P_{X},P_{X}P_{Z},(1-P_{X})P_{Z})$. Therefore, $s^{x}$ and $s^{z}$ are independent variables, and we have $J_{y}=0$ in the stat-mech model. Hence, the two flavors of Ising spins, $\{\sigma^{x}\}$ and $\{\sigma^{z}\}$, completely decouple except for the terms involving $s^{x,z}(T+\frac{1}{2})$ due to the tHH gates. We can further exchange all the spins $\sigma^{x}_{l}(t)\leftrightarrow\sigma^{z}_{\bar{l}}(t)$ for all $t\leq T$, after which the two spin flavors are completely decoupled. Since the disorder distributions of $s^{x,z}_{l}$ and $r^{x,z}_{p}$ are translation invariant, we can relabel their spatial locations $l$ and $p$ as $\bar{l}$ and $\bar{p}$ as well. Therefore, the only difference here, compared to the original stat-mech model in Eq. (C), is that the coupling constants $J_{x}$ and $J_{z}$, as well as $K_{x}$ and $K_{z}$, are exchanged for all the terms that involve the terms that are located within $t\leq T$. Consequently, both flavors of spins will locally confine once the error rates exceed the error threshold of quantum memory.
@@ C.2 Fold-transversal $S$ gate
Additionally, the toric code has a transversal $S$ (tSS) gate [53]. Physically, it can be implemented by first applying $S$ or $S^{\dagger}$ gates to every physical qubit on a diagonal, while also applying controlled-$Z$ (CZCZ) gates between pairs of qubits that are reflection-symmetric with respect to the diagonal line, see Fig. 12. For toric code on a torus with equal periodicity in both directions, the tSS gate simultaneously performs $S$ gates on the two encoded logical qubits.
@@
We now discuss the impact of the tSS gate on the error threshold using the stat-mech model. Again, we assume the independent-error model, in which $X$ and $Z$ errors occur independently, so that $J_{z}=0$ in the Hamiltonian. Therefore, the two flavors of Ising spins couple only through terms involving $s^{z}_{l}(T+\frac{1}{2})$. Since the disorder distributions are translation invariant, we can mirror-reflect the spatial location of all the $\sigma^{x}$ spins, so that the couplings are geometrically local. We recognize that, the stat-mech model in this case is essentially the same as Eq. (4) for two toric code blocks undergoing a tCNOT gate, where the two spins for the control and target code blocks become $\{\sigma^{x}\}$ and $\{\sigma^{z}\}$ spins for $X$ and $Z$ errors, and the bit-flip error rates of the two code blocks become the bit-flip and phase-flip error rate of the single code block. Therefore, for the tSS gate, the error threshold of logical $Z$ error will be locally lowered near $t=T$, similar to the target block of the tCNOT gate. In the stat-mech model, this effect can be effectively captured by the increase in local temperature of the $\sigma^{z}$ spins, as sketched in Fig. 10.
@@ Appendix D Stat-mech mapping for generic CSS codes under transversal Clifford gates
In this section, we generalize the stat-mech mapping approach to generic Calderbank-Steane-Shor (CSS) codes that support global transversal Clifford gates.
@@ D.1 Transversal logical circuit in the binary matrix formalism
The error-detection mechanism of a CSS code under repeated noisy syndrome extractions can also be understood using an extended version of the binary matrix formalism. Here, Pauli and syndrome errors over spacetime can be parameterized by a single binary vector $\mathfrak{e}$ over time [^4] In this section, the Fraktur alphabet always denotes the binary vector representation of a spacetime error, while the vector notation is used for binary vectors of length $n$, i.e. the physical errors.:
@@
Here $r_{x/z}(t)\in\{0,1\}^{n_{x/z}}$ represents the absence or existence of the syndrome extraction error at each $X/Z$ check at time $t$. The vertical lines in $\mathfrak{H}$ separate the columns that correspond to the physical Pauli error and syndrome extraction errors, and the convention will be used in subsequent discussions. Hence, the spacetime detectors formed by repeated syndrome measurements, whose parities are represented by a binary vector $\mathfrak{d}$, can be written in the following matrix form:
@@
where $\mathfrak{d}$ collects the parity of the spacetime detectors, $\mathds{1}_{n_{x/z}}$ is the identity matrix of dimension $n_{x/z}$, and $d^{x/z}(t+\frac{1}{2})\in\{0,1\}^{n_{x/z}}$ represents the parity of the $X/Z$ detectors. We will refer to $\mathfrak{H}$ as the spacetime parity check matrix.
@@
$U\Omega U^{T}=\Omega,\ \Omega\equiv\left(\begin{array}[]{cc}\mathds{1}_{n}&0\\ 0&\mathds{1}_{n}\end{array}\right).$ (53)
@@
where $V_{U}\in SL(n_{x}+n_{z},\mathbb{F}_{2})$ represents the change each individual stabilizer under the transversal gate. In the detector error model, $V$ determines the detector geometry across the transversal gate. Consider a transversal gate $U$ applied between $t=T$ and $t=T+1$ rounds of syndrome extraction. Assuming the physical Pauli errors do not persist across the transversal gate $U$, the spacetime parity check matrix can be written in the following block form:
@@
$\mathfrak{v}_{U}\mathfrak{H}|_{\bar{U}(T+\frac{1}{2})}=\mathfrak{H}\mathfrak{W}|_{\bar{U}(T+\frac{1}{2})},\ \mathfrak{v}_{U}=\text{diag}\left(\cdots,\mathds{1}_{n_{x}+n_{z}},V_{U},V_{U},\cdots\right),\ \mathfrak{W}|_{\bar{U}(T+\frac{1}{2})}=\text{diag}\left(\cdots,{\color[rgb]{0,0,1}\mathds{1}_{n_{x}+n_{z}}},U,{\color[rgb]{0,0,1}V_{U}},U,{\color[rgb]{0,0,1}V_{U}},\cdots\right).$ (56)
Here $\mathfrak{v}_{U}$ is a block-diagonal matrix with $V_{U}$ starting from the row block that represents the detectors between $t=T$ and $t=T+1$ (i.e. the row where $V_{U}^{-1}$ is at in Eq. (55)), and $\mathds{1}_{n_{x}+n_{z}}$ before that. $\mathfrak{W}|_{\bar{U}(T+\frac{1}{2})}$ is a block-diagonal matrix with alternating $U$ and $V_{U}$ blocks starting from the columns that represent the physical errors at $t=T+\frac{1}{2}$ (i.e. the column on the right of $V_{U}^{-1}$ in Eq. (55)).
@@
$\mathfrak{W}|_{\bar{U}(T+\frac{1}{2})\bar{U}^{\prime}(T^{\prime}+\frac{1}{2})}$ $=\text{diag}\left(\cdots,{\color[rgb]{0,0,1}\mathds{1}_{n_{x}+n_{z}}},U,{\color[rgb]{0,0,1}V_{U}},\cdots,UU^{\prime},{\color[rgb]{0,0,1}V_{U}V_{U^{\prime}}},\cdots\right)$
$=\text{diag}\left(\cdots,{\color[rgb]{0,0,1}\mathds{1}_{n_{x}+n_{z}}},U,{\color[rgb]{0,0,1}V_{U}},\cdots\right)\text{diag}\left(\cdots,{\color[rgb]{0,0,1}\mathds{1}_{n_{x}+n_{z}}},U^{\prime},{\color[rgb]{0,0,1}V_{U^{\prime}}},\cdots\right)=\mathfrak{W}|_{\bar{U}(T+\frac{1}{2})}\mathfrak{W}|_{\bar{U}^{\prime}(T^{\prime}+\frac{1}{2})}.$ (57)
@@ D.2 Stat-mech mapping via the binary matrix formalism
Using the binary matrix formalism, we can now discuss the stat-mech mapping with a transversal Clifford gate in the most general setting. Starting from the case of quantum memory. An error model in this case can be specified via the probability distribution
@@
Two spacetime errors $\mathfrak{e}$ and $\mathfrak{e}^{\prime}=\mathfrak{e}+\delta\mathfrak{e}$ trigger the same set of detectors if they differ by a cycle $\delta\mathfrak{e}\in\ker(\mathfrak{H})$. The trivial cycles, which do not lead to logical errors, are generated by the following four sets of bitstrings
@@
where $l_{z,x}$ are binary vectors that represent logical operators of the original CSS code [^5] It does not matter which time the logical error $l_{z,x}$ is located in $\mathfrak{l}_{x,z}$. In fact, using linear combinations of $\mathfrak{r}^{x,z}$, we can always shift it in time.. In this way, every cycle $\delta\mathfrak{e}$ can be parameterized in the following way:
@@
Putting everything together and using the definition in Eq. (2.1), the logical error rate of the decoder is
@@
The quenched disorder, $\mathfrak{e}$, follows the distribution $\mathrm{prob}(\mathfrak{e})$. For structured CSS codes (such as topological codes) and a distribution of errors $\mathrm{prob}(\mathfrak{e})$, the partition function can be brought into the Boltzmann form of a known Hamiltonian.
Now consider the case with a transversal logical gate. From Eq. (56), we see that every trivial cycle $\mathfrak{c}^{\prime}$ of the new spacetime parity check matrix $\mathfrak{H}^{\prime}$ with transversal gates is related to the kernel of $\mathfrak{H}$ via $\mathfrak{c}^{\prime}=\mathfrak{W}^{-1}\mathfrak{c}$, $\mathfrak{c}\in\ker(\mathfrak{H})$, where $\mathfrak{W}$ can be computed using Eq. (D.1). Therefore, the new partition function reads
@@
Nevertheless, the two partition functions differ only locally near $t=T$. In fact, away from the transversal gate (i.e. when $t\geq T+1$), it is easy to verify that the new cycles $\mathfrak{W}^{-1}\mathfrak{r}^{x,z}_{i}(t\geq T)$ are linear combinations of the old ones: